OpenStudy (anonymous):

A function is shown below: f(x) = (1.07)x What does the function represent? Exponential growth of 7% Exponential decay of 7% Exponential growth of 93% Exponential decay of 93%

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

I picked A, I just want to be sure. Because I think it might be C also

OpenStudy (anonymous):

Definitely A.

OpenStudy (phi):

If "x" is the exponent, you should write it (1.07)^x but yes, that is exponential growth of 7%

OpenStudy (phi):

growth of 93% means the if you start with 100, after 1 step (x goes up by 1) you expect 93% more... 93% of 100 is 93, so 100+93 = 193 7% growth would be 100 + 7% of 100 = 100+7= 107

OpenStudy (phi):

at x=0 f(0)= 1 (because 1.07^0 is 1) at x=1 f(1) = 1.07 1.07 grew by 0.07 which is 7% of 1

OpenStudy (anonymous):

Ohh